The resistivity of gold is 2.44 x 10-8 ohms.m at room temperature. A gold wire that is 0.9 mm in diameter and 14 cm long carries a current of 940 mA. What is the electric field in the wire?
Answer

0.090 V/m
0.028 V/m
0.046 V/m
0.0090 V/m
0.036 V/m
Answer #1
Concepts and reason

The concept needed to solve this problem is the expression for electric field inside a current carrying conductor.

Write the expression for electric field inside a current conductor. Substitute all the values in this equation and solve for required value of electric field.

Fundamentals

The expression for electric field (E) inside a current conductor is,

E=ρJE = \rho J

Here, J is the current density and ρ\rho is the resistivity of the conductor.

The current density is defined as the current divided by cross section area of the conductor.

J=IAJ = \frac{I}{A}

The cross section area of a wire of radius r can be calculated using the following formula:

A=πr2A = \pi {r^2}

The relation between the radius and diameter of a circle is,

r=d2r = \frac{d}{2}

The cross section area of a wire of radius r can be calculated using the following formula:

A=πr2A = \pi {r^2}

The relation between the radius and diameter of a circle is,

r=d2r = \frac{d}{2}

Substitute d2\frac{d}{2} for r in the equation A=πr2A = \pi {r^2} .

A=π(d2)2=πd24\begin{array}{c}\\A = \pi {\left( {\frac{d}{2}} \right)^2}\\\\ = \frac{{\pi {d^2}}}{4}\\\end{array}

Substitute πd24\frac{{\pi {d^2}}}{4} for A in the equation J=IAJ = \frac{I}{A} .

J=I(πd24)=4Iπd2\begin{array}{c}\\J = \frac{I}{{\left( {\frac{{\pi {d^2}}}{4}} \right)}}\\\\ = \frac{{4I}}{{\pi {d^2}}}\\\end{array}

Convert the unit of diameter from millimeter to meter:

d=0.9mm(103m1mm)=0.9×103m\begin{array}{c}\\d = 0.9{\rm{ mm}}\left( {\frac{{{{10}^{ - 3}}{\rm{ m}}}}{{1{\rm{ mm}}}}} \right)\\\\ = 0.9 \times {10^{ - 3}}{\rm{ m}}\\\end{array}

Convert unit of current from mA to A:

I=940mA(103A1mA)I = 940{\rm{ mA}}\left( {\frac{{{{10}^{ - 3}}{\rm{ A}}}}{{1{\rm{ mA}}}}} \right)

Substitute 0.9×103m0.9 \times {10^{ - 3}}{\rm{ m}} for d and 940×103A940 \times {10^{ - 3}}{\rm{ A}} for I.

J=4(940×103A)π(0.9×103m)2=1.478×106A/m2\begin{array}{c}\\J = \frac{{4\left( {940 \times {{10}^{ - 3}}{\rm{ A}}} \right)}}{{\pi {{\left( {0.9 \times {{10}^{ - 3}}{\rm{ m}}} \right)}^2}}}\\\\ = {\rm{1}}{\rm{.478}} \times {10^6}{\rm{ A/}}{{\rm{m}}^2}\\\end{array}

The expression for electric field inside a current conductor is,

E=ρJE = \rho J

Substitute 1.478×106A/m2{\rm{1}}{\rm{.478}} \times {10^6}{\rm{ A/}}{{\rm{m}}^2} for J and 2.44×108Ω2.44 \times {10^{ - 8}}{\rm{ }}\Omega for ρ\rho .

E=(1.478×106A/m2)(2.44×108Ω)=0.0360632V/m=0.036V/m\begin{array}{c}\\E = \left( {{\rm{1}}{\rm{.478}} \times {{10}^6}{\rm{ A/}}{{\rm{m}}^2}} \right)\left( {2.44 \times {{10}^{ - 8}}{\rm{ }}\Omega } \right)\\\\ = {\rm{0}}{\rm{.0360632 V/m}}\\\\ = {\rm{0}}{\rm{.036 V/m}}\\\end{array}

Ans:

The value of electric field inside the wire is 0.036V/m{\rm{0}}{\rm{.036 V/m}} .